Solve the following linear differential equations

1) dy/dx + 2y = sinx

2) dy/dx + y = e^x

3) dy/dx +(4/x)y = x³y³




1
Expert's answer
2021-01-26T03:21:08-0500

1) "\frac{dy}{dx}+2y=\sin \left(x\right)"  - First order linear Ordinary Differential Equation

A first order linear ODE has the form of     "y'\left(x\right)+p\left(x\right)y=q\left(x\right)"  

Substitute   "\frac{dy}{dx}\mathrm{\:with\:}y'\:"

"y'\:+2y=\sin \left(x\right)"

The equation is in the first order linear ODE form

We find the integration factor:  "\mu(x) = e^{2x}"

We put the equation in the form "(\mu(x)y)' = \mu(x)q(x):"     "(e^{2x}y)' = sin(x)e^{2x}"

Solve "(e^{2x}y)' = sin(x)e^{2x} :"            "y=-\frac{\cos \left(x\right)}{5}+\frac{2\sin \left(x\right)}{5}+\frac{c_1}{e^{2x}}"

Answer: "y=\large-\frac{\cos \left(x\right)}{5}+\frac{2\sin \left(x\right)}{5}+\frac{c_1}{e^{2x}}"


2) "\frac{dy}{dx}+y=e^x" -  First order linear Ordinary Differential Equation

A first order linear ODE has the form of    "y'\left(x\right)+p\left(x\right)y=q\left(x\right)"

Substitute  "\frac{dy}{dx}\mathrm{\:with\:}y'\:"

"\frac{dy}{dx}+y=e^x"

The equation is in the first order linear ODE form

We find the integration factor: "\mu(x) = e^{2x}"

We put the equation in the form "(\mu(x)y)' = \mu(x)q(x):"   "(e^xy)'=e^{2x}"

Solve "(e^xy)'=e^{2x}:"   "y=\frac{e^x}{2}+\frac{c_1}{e^x}"

Answer: "y=\large\frac{e^x}{2}+\frac{c_1}{e^x}"


3) "\frac{dy}{dx}+\left(\frac{4}{x}\right)y=x^3y^3" -   First order Bernoulli Ordinary Differential Equation

A first order Bernoulli ODE has the form of    "y'+p\left(x\right)y=q\left(x\right)y^n"

Substitute  "\frac{dy}{dx}\mathrm{\:with\:}y'\:"

"y' + \left(\frac{4}{x}\right)y=x^3y^3"

The equation is in first order Bernoulli ODE form

The general solution is obtained by substituting "\nu = y^{1-n}" and solving "\frac{1}{1-n}\nu' + p(x)\nu=q(x)"

Transform to  "\frac{1}{1-n}\nu' + p(x)\nu=q(x)" :    "\large-\frac{\nu'}{2}+\frac{4\nu}{x}=x^3"

Solve "\large-\frac{\nu'}{2}+\frac{4\nu}{x}=x^3" :   "\nu = \large\frac{x^4}{2}+c_1x^8"

Substitute back "\nu = y^{-2}:"      "y^{-2} = \large\frac{x^4}{2}+c_1x^8"

Isolate y:     "\large y=\sqrt{\frac{2}{x^4\left(1+2c_1x^4\right)}},\:y=-\sqrt{\frac{2}{x^4\left(1+2c_1x^4\right)}}"

Answer:  "\large y=\sqrt{\frac{2}{x^4\left(1+2c_1x^4\right)}},\:y=-\sqrt{\frac{2}{x^4\left(1+2c_1x^4\right)}}"